# Tailoring the Hyperparameters of a Wide-Kernel Convolutional Neural Network to Fit Different Bearing Fault Vibration Datasets

### v1

@@ Abstract
State-of-the-art algorithms are reported to be almost perfect at distinguishing the vibrations arising from healthy and damaged machine bearings, according to benchmark datasets at least. However, what about their application to new data? In this paper, we are able to confirm that neural networks for bearing fault detection can be crippled by incorrect hyperparameterisation, and also that the correct hyperparameter settings can actually change when transitioning to new data. The paper weaves together multiple methods to explain the behaviour of the hyperparameters of a wide-kernel convolutional neural network and how to set them. Since guidance already exists for generic hyperparameters like minibatch size, we focus on how to set architecture-specific hyperparameters such as the width of the convolutional kernels, a topic which might otherwise be obscure. We reflect different data properties by fusing information from seven different benchmark datasets, and our results show that the kernel size in the first layer in particular is sensitive to changes in the data. Looking deeper, we use manipulated copies of one dataset in an attempt to spot why the kernel size sometimes needs to change. The relevance of sampling rate is studied by using different levels of resampling, and spectral content is studied by increasingly filtering out high frequencies. At the end of our paper we conclude by stating clear guidance on how to set the hyperparameters of our neural network architecture.
keywords:
Deep Learning , CNN , Hyperparameter Analysis , Time Series Classification , Industrial Fault Detection , Bearing Fault Detection
journal: the arXiv preprint server
@@
[aff:osnabrueck]organization=Semantic Information Systems Group, Osnabrück University,addressline=Wachsbleiche 27, postcode=49069, city=Osnabrück, state=Lower Saxony, country=Germany
[…]
[aff:dfki]organization=German Research Center for Artificial Intelligence (DFKI),
addressline=Hamburger Straße 24, postcode=49084, city=Osnabrück, state=Lower Saxony, country=Germany
1 Introduction
@@
Machines are full of bearings, the components that make rotation possible. They are small metal balls usually arranged in a circular casing and by rolling, these spheres allow one piece of a machine to move around relative to another piece of the machine. Wear and tear unfortunately can accumulate over time into damage that causes the bearing element to ultimately break down, taking the entire machine offline with it. For example, one way this can happen is if small scratches develop in the casing that holds the bearings in place. As the ball bearings roll over that scratch again and again (perhaps hundreds of times per minute for days, weeks and months), it can grow into a structural weakness. One day, the casing can crack due to this weakness, breaking the rotating element and putting the entire machine out of commission until repairs are made. Therefore, methods for detecting such faults are crucial.
@@
In this paper – a significantly extended and adapted revision of [1] building on our work presented in [2] – we tackle the question of how to configure neural networks so that they can effectively identify different types of damaged bearings. Thus, in this paper we do not start from scratch but instead our work takes place after and continues from research we published earlier. In particular, this paper presents a series of new experiments that address questions raised and left unresolved by our earlier work in [1] and [2]. Two new neural network architectures are studied in order to show how important hyperparameter configuration is to a wider range of networks. We present a novel experimentation methodology, extending the methodology presented in [2], for assessing the impact of hyperparameters in the context of wide-kernel approaches for bearing fault detection. Doing that, we run new experiments on manipulated subsets of the data to see how dataset properties affect the key question of how to configure hyperparameters. Specifically, we present more datasets, more neural network architectures and more extensive results compared to [2].
1.1 Finding Faults and Beating Benchmarks
From outside of the machine, the damage that develops in a bearing element cannot be seen. Bearings are encased within the machine and hidden from view. This creates obvious problems for monitoring the accumulation of damage within a bearing element and deciding when a replacement is needed. However, using vibrations which spread through the machine, we can try to look inside at what is happening: sensors placed on the outside of the machine can pick up vibrations that originate from deep inside. This is an approach that gives some hope of detecting when replacements are needed before a machine has an unexpected breakdown. Unfortunately, vibrations bring their own problems since reading these signals is not like looking at a photograph, and instead a lot of work is needed to interpret the vibrations that are being gathered by the sensors. This is in particular relevant to initiatives such as Industry 4.0 [3, 4], e. g., for enabling predictive maintenance applications [5], for which also explainable and interpretable approaches are usually beneficial [6].
One way to interpret vibration data and figure out how to distinguish healthy and damaged bearings is to learn from experience, by applying machine learning methods. Neural networks in particular are machine learning methods that can be very successfully trained by exposing them to a dataset of examples. As summarised in [7], various types of neural networks have been reported to achieve 97%+ accuracy on fault detection tasks, in papers such as [8, 9, 10, 11]. Furthermore, we also developed various deep learning algorithms to improve performance in fault detection, as described in [12, 13].
In the pursuit of effective algorithms to recognise when damage has occurred to a bearing element, researchers have settled on a series of ‘benchmark’ datasets, used to train and test different approaches. The obvious benefit of these openly-available datasets is that it is like having a fair race: everyone is starting at the same place and running the same route, so we can compare algorithms fairly based on their performance on these benchmark datasets. In our work, we make use of seven datasets originating from different organisations: Case Western Reserve University (CWRU), Paderborn University, the so-called ‘Gearbox’ dataset, the Society for Machine Failure Prevention Technology (MFPT), University of Conneticut (UoC), Southeast University (SEU) and the Xi’an Jiaotong University (XJTU).
@@ 1.1 Finding Faults and Beating Benchmarks
Looking at past results on benchmark datasets, it has become obvious that multiple successful deep learning approaches exist. A recent review of the literature on the topic [7] indicates that at least 25 different deep learning algorithms have achieved over 95% accuracy (often 99% or 100%) on one or more of the benchmark datasets mentioned above. Such consistently high levels of performance imply that the difference between two state-of-the-art algorithms is now predestined to be small and there is little room left to improve beyond previous results. Therefore, comparing one state-of-the-art option to another is becoming less interesting and important than before.
Previous research in bearing fault detection has focused almost exclusively on finding algorithmic innovations with the hope of achieving higher scores during testing; however, since there is little room to improve, we now propose that the salient question is not whether new algorithms can be found, but whether the existing algorithms will continue to perform well when applied to new, unseen data. This is the question of whether they will still achieve near-perfect results when someone applies them in a real-life setting, when the sensors and machines are different from what the benchmarks used.
1.2 Limits of What the Benchmarks Can Tell Us
Benchmarks are very useful when comparing algorithms but they can only give an indication about how successful a method will be when applied to new data in the real world. The belief that an algorithm will perform well when it analyses a new set of machine vibrations, in a previously-unseen context, is simply an assumption and an assumption that can be challenged. For example, the types of damage that build up when a machine is really used in a factory might go beyond what is seen amongst the artificially-created damage recorded in benchmark datasets. Who can say that the types of damage found in the real factory will not be more subtle and hard to detect, or more diverse and thus requiring an algorithm with increased flexibility? Such questions start to introduce uncertainty. Differences in the properties of the recorded data might mean that an algorithmic approach does not generalise perfectly, leading to inferior performance when it is applied to a new dataset.
Speaking more generally about neural networks, when a practitioner takes an existing algorithm and tries using it on their own data, there is a risk that it does not work. Imagine that the practitioner is left with an algorithm that does not work. What can they do? One of the first things they can try is to modify the hyperparameters of the algorithm. In order to control the training of a neural network, the user must specify some so-called ‘hyperparameters’, such as the minibatch size (the number of examples used in a single step of training). These hyperparameters control how to generate the training signal that the neural network learns from – they are used in generating the numbers that incrementally update the neural network. Just like how a mis-tuned radio will not play your favourite radio show, it is possible to mis-tune hyperparameters and end up with a training signal that does not do what you want, i.e., it will not teach the neural network how to do the task.
1.3 The Open Question of Hyperparameters
Alongside hyperparameters that modulate the training process, one can consider what we might call ‘architectural hyperparameters’, which control the size and shape of a neural network. Quite commonly, when faced with poor results, a practitioner might make their neural network bigger. The layers used to construct the network could be enlarged or the network could be built by stacking more layers together. After making such changes, the neural network will contain more trainable parameters (numbers used to store information about how to process inputs) and thus have more capacity to learn about the task. If a neural network is not working then quite a natural step is to make it larger by increasing the architectural hyperparameters.
The reason we tweak hyperparameters is that the wrong choice of settings will lead to inferior performance whilst good choices will bring out the best from a neural network. In the case of severe ill-hyperparameterisation, a network might fail to learn at all during training. Returning specifically to bearing fault detection, we noted earlier that the differences appear to be minimal between the top 25 neural networks for detecting damaged bearings. By contrast, larger differences in accuracy can emerge when the hyperparameters of a network are changed. For example, [14] reports that the accuracy of a bearing fault classifying CNN is cut in half simply as a result of increasing from 6 convolutional layers to 8. When it comes to getting near-perfect accuracy on bearing fault detection, getting the hyperparameters right might be more important than choosing between neural network architectures.
However, although bad hyperparameter settings can result in misery, there is a question of why this matters, because we know that the hyperparameters work on our benchmark data. If the hyperparameters are already configured well, then what exactly is the problem? This is really a question of whether hyperparameters that work well on one dataset will necessarily be successful on another. To hone in on just a single reason why this might not always happen, we can ask whether drastically changing the sampling rate would impact what size the filters in a convolutional network should be.
Convolutional filters act like learned patterns that the network recognises at different time points during the input signal. Each pattern is stored in an array of numbers. If we increase the sampling rate used to record a signal that is otherwise kept fixed – so that more points are used to record the same peaks, troughs and plateaus over time – then it might be reasonable to expect that the number of points used to record the important patterns (in convolutional filters) also needs to be increased. The main point of this example is to highlight how data properties like the sampling rate might cause a meaningful difference in what hyperparameters are successful on different datasets. In fact, by examining the performance of neural networks on seven benchmark datasets, one important contribution of our work will be to show that hyperparameters working on one dataset sometimes stop working when changing to another dataset.
Another question is whether some hyperparameters can be changed without affecting the accuracy of the neural network on a specific dataset. Of course, it is possible to imagine that there is a miniscule ‘sweet spot’ of hyperparameters, a specific combination of numbers which is the only one that works. The reason this seems unlikely is that there is no obvious explanation for how the authors of successful algorithms were able to find this one-and-only sweet spot. Therefore, we instead expect that at least some of the hyperparameters can be modified without disastrous consequences for the accuracy of the neural network. Having more options in this way, it would be easier for an author to find a successful combination of hyperparameters without performing a rigorous search.
1.4 Shaping the Neural Network
@@
Architectural hyperparameters like the number of layers in a network affect how complicated it is to process each input. Making the network smaller reduces the number of internal parameters it has, meaning that the computer needs to perform fewer arithmetic operations. This has a knock-on effect that impacts, for example, how long it takes to train the network. Smaller networks train faster and are quicker to generate predictions on new data. If some architectural hyperparameters can be reduced without harming the accuracy, it would be good to know this. Practitioners would benefit from knowing that they can make the network smaller and thus faster to train and use. Moreover, in real-world settings, there is the possibility of ‘embedding’ fault detection algorithms, so that they run on small devices situated in and amongst the machines they analyse rather than running on a user’s laptop. Embedded processing skips the step of extracting data from the sensors and communicating it to a central location (like a laptop), and removes the need to keep a laptop or computer running continuously in order to process the data. An additional benefit of small networks is that they can be run on smaller and cheaper embedded devices. Clearly, there is real value to understanding the hyperparameters of a network, to see which hyperparameters can and cannot be changed when processing a given dataset. In this vein, we have worked to understand the architectural hyperparameters of a network capable of state-of-the-art results on fault detection tasks.
Of course, there must be limits to how much we can shrink a network or otherwise manipulate its hyperparameters. Some changes will be too drastic to work and they will make the neural network less able to detect bearing faults. Another line of our research is to study which hyperparameters are crucial and require the most careful handling. We wish to understand how much impact different changes will have and be able to state the relative importance of the hyperparameters. Re-examining the wide-kernel network presented in  [15, 12], we distil new explanations of how the network’s hyperparameters impact its accuracy at classifying bearing fault signals.
1.5 Do Dataset Properties Determine Performance?
Datasets have different properties, one reason why hyperparameters might need (re)tuning on new data even though they were already working on benchmarks. In the field of bearing vibration, research focuses on sensor recordings stored in ‘time series’, so called because they consist of a series of sequential measurements of some variable over time. The sampling rate is one of the key data properties when discussing time series. With a sampling rate of 12kHz, the movement of a vibrating machine is measured every 0.00008333 seconds. The sampling rate determines how many measurements are made per second and conversely how far apart in time each individual measurement point is. The sampling rate is documented for most benchmark datasets and we can see that it varies. One thread of our research will be to find out if changes to the sampling rate have an impact on the accuracy of a network and on which hyperparameter settings are best.
@@ 1.5 Do Dataset Properties Determine Performance?
The sampling rate affects the dataset in multiple ways. It affects what frequency bands can be detected, with half the sampling rate (also known as the ‘Nyquist frequency’) providing the upper limit of what frequencies can be captured. An increased sampling rate also means that more numbers are used to record a vibrating machine, implying that the neural network needs to do more computation in order to process the inputs. One open question is how much the highest frequencies in vibration recordings are used by neural networks. This is interesting because if the highest frequencies are not needed, then storing and processing the recordings could be made more efficient by reducing the sampling rate. In fact, if high frequencies are unnecessary for fault detection then removing them could be a form of noise filtering. This is not implausible, since many signals resulting from vibrations (like from musical instruments) have an overtone series, implying that there is redundancy. As noted in [16], vibrations in bearing elements are transmitted to the sensors through the rest of the machine, which also consists of multiple components that themselves vibrate. This potentially leads to irrelevant or misleading information in the recorded signal, and if some frequency bands are just noise then it could be beneficial to remove them. Our research looks at how changes in the sampling frequency, for example when moving to one dataset to another one, affect neural networks. We ask whether filtering the irrelevant frequencies be helpful, for example by increasing the size of the ‘sweet spot’ of hyperparameters.
1.6 Explaining Hyperparameters
How many layers is enough in a neural network? Why do we set other hyperparameters in the way that we do? Although there are rules to follow when tuning some hyperparameters like the minibatch size (e.g., “set it to the largest value that will still allow the minibatch to fit into GPU memory”) in other cases we can look at a neural network and find that the hyperparameter settings seem totally arbitrary. To pick just a pair of fault detection papers at random, we find that [9] uses 20 hidden units in fully-connected layers whilst [17] uses 64 instead. As is typical in the field, neither author noted why they chose the particular numbers they use and the reader is left with no explanation of why there is a difference. If there is ever a need to adapt these hyperparameters when working with new datasets, future users could quite justifiably feel lost. What users in fact want is a simple, explainable process for setting the hyperparameters.
The terms “Explainability” and “Explainable AI (XAI)” tie together several strands of research aimed at clarifying how the inputs to an algorithm relate to the outputs. With complex statistical models such as neural networks the inner workings are not understandable and we must rely upon specialised techniques to illuminate why particular outputs were chosen by the algorithm. This is a large topic and there are many surveys of the work done in this research area, such as [18] for example. What XAI aims to do is to cast a light on obscure and hidden relationships that exist in the algorithm and which link the inputs to the outputs. In many real-world domains such as bearing fault detection, algorithms need to learn nonlinear mathematical functions in order to approximate the complex interactions between multiple factors seen in observed data.
Returning now to our own focus, we want to do something similar with hyperparameters in order to better understand how they propel some networks to near-perfect accuracy and other networks to near-total failure. As we have noted, neural networks are clearly powerful and numerous in bearing fault classification but they are also poorly understood. By adapting techniques from XAI, we aim to build a bridge between empirical results and more general insights about how to tune hyperparameters.
1.7 The Paper Ahead
@@
The present paper re-examines and then extends upon the results from several pieces of our own past work. In addition to providing new insights on past results, this present study presents new experiments with resampling and filtering of the data, allowing us to investigate how the properties of the data affect hyperparameter tuning.
To summarise the intention of the research: we already have neural network architectures for bearing fault detection ([7] gives many examples) but their success relies on hyperparameters which are poorly understood. Hyperparameters become especially important when we remember that the users of a neural network will not want to take it and apply it to a benchmark dataset. For the network to be any use at all, it must be possible for someone who is not the developer to take it and apply it themselves to novel datasets. To do so, they need some idea of how to decide the hyperparameter settings. However, the hyperparameters might as well be written in a cypher – there is no intuitive way to know that 64 convolutional filters works but a rate of 128 does not. The numbers seem arbitrary.
Engineers in a factory might have different machinery and types of bearings from what was used in benchmark datasets. They may have restrictions about where the sensors can be placed, and even what sensors they can use to record vibrations. Together, the specifics of a factory’s machinery, bearing types, sensor availability and sensor placement will conspire to make a unique dataset for each engineer. As already noted, the choice of algorithm will not impact performance that much on benchmark datasets (see [7]), but the performance drop as a result of trying out a new dataset might be noticeable. Therefore, we take an existing neural network architecture which is capable of achieving state-of-the-art results on multiple benchmarks, and enrich it by presenting a detailed analysis of the hyperparameter decisions presented to the user.
@@
We use a cocktail of multiple methods to explain the impact of hyperparameters on the wide-kernel CNN’s accuracy. We summarise results from past work in order to see the average performance level when specific hyperparameter values are carried across seven datasets. We also harvest new results from experiments on manipulated versions of the data. We quantify the ‘feature importance’ – the extent to which different variables contribute to an outcome – of different hyperparameters and how it changes when processing either resampled or filtered copies of a dataset. This lets us find out how relevant dataset properties like the sampling rate are to deciding what hyperparameters to use. We also undertake separate additional experiments with two other architectures (besides the wide-kernel CNN) in order to round out the generality of our basic premise.
@@
  • 2.
    We show that hyperparameters affect how accurate a neural network is (the majority of our work specifically investigates a wide-kernel CNN architecture).
  • 3.
    Fusing information from 7 datasets to analyse the wide-kernel architecture, we show that hyperparameters that are successful on one dataset are not necessarily successful on another dataset.
  • 4.
    We also highlight that not all hyperparameters are equally important.
  • 5.
@@
    After fusing the information from multiple different experiments, we find that due to redundancy in the signal, very high frequencies are not needed for successful bearing fault detection.
  • 7.
@@
    We give general guidelines for how to set the hyperparameters of a wide-kernel CNN when transitioning to new data.
2 Background & Related Work

### v2

@@ Abstract
State-of-the-art algorithms are reported to be almost perfect at distinguishing the vibrations arising from healthy and damaged machine bearings, according to benchmark datasets at least. However, what about their application to new data? In this paper, we confirm that neural networks for bearing fault detection can be crippled by incorrect hyperparameterisation, and also that the correct hyperparameter settings can change when transitioning to new data. The paper combines multiple methods to explain the behaviour of the hyperparameters of a wide-kernel convolutional neural network and how to set them. Since guidance already exists for generic hyperparameters like minibatch size, we focus on how to set architecture-specific hyperparameters such as the width of the convolutional kernels, a topic which might otherwise be obscure. We reflect different data properties by fusing information from seven different benchmark datasets, and our results show that the kernel size in the first layer in particular is sensitive to changes in the data. Looking deeper, we use manipulated copies of one dataset in an attempt to spot why the kernel size sometimes needs to change. The relevance of sampling rate is studied by using different levels of resampling, and spectral content is studied by increasingly filtering out high frequencies. We find that, contrary to speculation in earlier work, high-frequency noise is not the main reason why a wide kernel is preferable to a narrow kernel. Finally, we conclude by stating clear guidance on how to set the hyperparameters of our neural network architecture to work effectively on new data.
Index Terms:
Bearing Fault Detection, CNN, Deep Learning, Industrial Fault Detection, Hyperparameter Analysis, Time Series Classification
@@
=-15pt
I Introduction
@@
Machines are full of bearings, small metal balls usually arranged in a circular casing, which roll around and allow one piece of a machine to rotate relative to another. Even during their ordinary working life it is possible for bearing components to pick up damage such as, e.g., a tiny scratch in the casing. As the ball bearings roll over that scratch again and again (perhaps hundreds of times per minute for days, weeks and months), it can grow into a structural weakness. One day, the casing can crack due to this weakness, breaking the rotating element and putting the entire machine out of commission until repairs are made. Therefore, methods for detecting such faults are crucial. Unfortunately the accumulating damage cannot be seen from outside the machine, and so practitioners often turn to recorded vibrations which spread easily from the inside to the outside.
At the time of writing, the most effective method for analysing vibration signals in order to identify faulty bearings is to use neural networks. Neural networks have demonstrated excellent accuracy at classifying different types of bearing vibrations but their high accuracy is not guaranteed. Even the best-performing neural networks have multiple ‘hyperparameters’ – decisions that are left to the user yet which can determine whether the network succeeds or not. We contend that although hyperparameters are crucial they are under-researched within the field of bearing fault research. Throughout the present paper we take a recent state-of-the-art neural network and ask how its performance changes as a result of changing its hyperparameters. By the end of the paper we can sketch a map of the “hyperparameter landscape” of our chosen network, showing how the neural network responds as we traverse combinations of hyperparameters.
@@
This paper is a significantly extended and adapted revision of [1] building on our work presented in [2]. We present a series of new experiments that address questions raised and left unresolved by our earlier work in [1] and [2]. Two new neural network architectures are studied in order to show how important hyperparameter configuration is to a wider range of networks for bearing fault detection. Presenting a novel experimentation methodology, extending the methodology presented in [2], we run new experiments on manipulated subsets of the data to see how two main dataset properties (sampling rate and spectral content) affect the hyperparameter landscape. Overall, we present more datasets, more neural network architectures and more types of experiments compared to [2]. In the next subsections of this introduction, we briefly lay out the rationale of our work.
I-A Finding Faults and Beating Benchmarks
In the pursuit of effective algorithms to recognise when damage has occurred to a bearing element, researchers have settled on a series of ‘benchmark’ datasets, used to train and test different approaches. These openly-available datasets act like a fair contest through which we can compare the accuracy of different algorithms. Operating on well-known benchmark data, various types of neural networks have been reported to achieve 97%+ accuracy on fault detection tasks, in papers such as [3, 4, 5, 6]. A recent review of the literature on the topic [7] indicates that at least 25 different deep learning algorithms have achieved over 95% accuracy (often 99% or 100%) on one or more bearing fault benchmarks.
@@ I-A Finding Faults and Beating Benchmarks
Such consistently high levels of performance imply that the difference between two state-of-the-art algorithms is now predestined to be small and there is little room left to improve beyond previous results. Therefore, comparing one state-of-the-art option to another is becoming less interesting and important than before. What we find more interesting is the question of whether they will still achieve near-perfect results when someone applies them in real life, when the sensors and machines are different from what the benchmarks used.
Imagine that a practitioner tries an existing algorithm on their own data and it does not work. What can they do? One of the first things they can try is to modify the so-called ‘hyperparameters’, decisions about how to structure and train the network. It has been noted that the performance of most machine learning algorithms depends a lot on hyperparameters. Tuning hyperparameters to their best values can be at least as important as the choice of algorithm [8] or chasing after marginal gains on the state-of-the-art by developing new architectures [9]. For example, [10] reports that the accuracy of a bearing fault classifying convolutional neural network is cut in half simply as a result of increasing from 6 convolutional layers to 8. When it comes to getting near-perfect accuracy on bearing fault detection, getting the hyperparameters right might be more important than choosing between neural network architectures.
I-B Shaping the Neural Network
@@
Architectural hyperparameters like the number of layers in a network affect how complicated it is to process each input. Making the network smaller reduces the number of internal parameters it has, meaning that the computer needs to perform fewer arithmetic operations. This has a knock-on effect meaning that smaller networks train faster and are quicker to generate predictions on new data. If some architectural hyperparameters can be reduced without harming the accuracy, it would be good to know this. Practitioners would benefit from knowing that they can make the network smaller and thus faster to train and use. Moreover, in real-world settings, there is the possibility of ‘embedding’ fault detection algorithms, so that they run on small devices situated in and amongst the machines they analyse rather than running on a user’s laptop. This would require networks that can run using reduced storage, memory and processor resources.
Of course, there must be limits to how much we can shrink a network or otherwise manipulate its hyperparameters. Some changes will be too drastic to work and they will make the neural network less able to detect bearing faults. How many layers is enough in a neural network? Why do we set other hyperparameters in the way that we do? Picking two fault detection papers to use as an example, we find that [4] uses 20 hidden units in fully-connected layers whilst [11] uses 64 instead. As is typical in the field there is no stated reason for why these particular numbers should be used, resulting in the danger that if future users ever need to adapt these hyperparameters for new datasets, they could justifiably feel lost. We suspect that what users in fact want is not just an algorithm that performs well on benchmarks, but also a simple process for setting the hyperparameters when moving to new data. We use an approach called “multiple defaults” to very efficiently tune the hyperparameters of a neural network for bearing fault classification.
I-C Do Dataset Properties Determine Performance?
Datasets have different properties, providing one reason why hyperparameters might need (re)tuning on new data even though they were already working on benchmarks. To record vibrations as time series leads to them having several important properties, including the sampling rate which determines how many measurements are made per second and conversely how far apart in time each measurement point is. Sampling at 12kHz, the movement of a vibrating machine is measured every 0.00008333 seconds. The sampling rate affects the dataset in multiple ways. First, an increased sampling rate means that more numbers are used to record a vibrating machine, implying that the neural network needs to do more computation in order to process the inputs. Likewise it is plausible that the size of ‘convolutional filters’, which act like stored patterns in certain neural networks, needs to be bigger when dealing with a higher sampling rate.
@@ I-C Do Dataset Properties Determine Performance?
The sampling rate additionally affects what frequency bands can be detected, with half the sampling rate (also known as the ‘Nyquist frequency’) providing the upper limit of what frequencies can be captured. One open question is how much the highest frequencies in vibration recordings are used by neural networks. This is interesting because if the highest frequencies are not needed, then storing and processing the recordings could be made more efficient by reducing the sampling rate. In fact, if high frequencies are unnecessary for fault detection then removing them could be a form of noise filtering. This is not implausible, since many signals resulting from vibrations (like from musical instruments) have an overtone series, implying that there is redundancy. As noted in [12], vibrations in bearing elements are transmitted to the sensors through the rest of the machine, which also consists of multiple components that themselves vibrate. This potentially leads to irrelevant or misleading information in the recorded signal, and if some frequency bands are just noise then it could be beneficial to remove them. Our research looks at how a wide-kernel convolutional network responds to changes in the sampling frequency. To further pinpoint the matter, we also ask whether filtering the irrelevant frequencies without changing the sampling rate is helpful. This leads us to manipulate one of the benchmark datasets in two ways: resampling, and, filtering.
I-D The Paper Ahead
@@
The paper ahead re-examines and then extends upon the results from several pieces of our own past work. In addition to providing new insights on past results, this present study presents new experiments with resampling and filtering of the data, allowing us to investigate how the properties of the data affect hyperparameter tuning.
@@
We use a cocktail of multiple methods to explain the impact of hyperparameters on a wide-kernel CNN’s accuracy. We summarise results from past work in order to see the average performance level when specific hyperparameter values are carried across seven datasets. We also gather new results from experiments on manipulated versions of the data. This lets us find out how relevant dataset properties like the sampling rate are to deciding what hyperparameters to use. By specifying multiple defaults – a sequence of hyperparameter settings to try out on new data – we give users robust guidance on how to tune our chosen wide-kernel network architecture on new data. We also undertake separate additional experiments with two other architectures (besides the wide-kernel CNN) in order to round out the generality of our basic premise.
@@
  • After integrating the information from multiple different experiments, we find that due to redundancy in the signal, very high frequencies are not needed for successful bearing fault detection.
@@
  • We give a precise process, based on the notion of “multiple defaults”, for how to set the hyperparameters of a wide-kernel CNN when transitioning to new data.
II Background & Related Work